OpenStudy (anonymous):

Find the exact value of the following function. Show your work and explain the steps you take to find the value. sec(17π/6)

OpenStudy (anonymous):

use the fact that 17/6=(10+7)/6

OpenStudy (anonymous):

sec a=1/cosa= so we shld get cos 17pi/6 \[\cos (A+B)=\cos Acos B-\sin Asin B\] \[A=10/6\pi\]\[B=7/6\pi\]

OpenStudy (anonymous):

Why was this bumped?